OpenStudy (anonymous):

How do you solve: If 3x/2 +2 = x/6, the x equals?

OpenStudy (anonymous):

x = -3/2

OpenStudy (anonymous):

3x/2 - x/6 = -2 9x/6 - x/6=-2 8x/6=-2 4x/3=-2 4x=-6 x=-6/4 x=-3/2